OpenStudy (anonymous):

prove THAT sin3x=cos3x

sam (.sam.):

Start by \[\sin(3x)=\sin(2x+x)\] Then use the sum to difference formula

OpenStudy (anonymous):

are there any values setfor this question? sin 3x = cos 3x only when tan 3x =1

OpenStudy (anonymous):

but not a complete answer

OpenStudy (anonymous):

wrong SHADOWYS

OpenStudy (anonymous):

oh. okay. sub x=0 into the eq and see what u get

OpenStudy (anonymous):

X is a angle not a variable

OpenStudy (anonymous):

In this case, since the question is asking for the [[proof]], not the solution, i don't see how you can get it, without setting the angles into set angles, like \(\frac{\pi}{12}\) rad

OpenStudy (anonymous):

i can give you a hint..!

OpenStudy (anonymous):

in the case where you need the solution, sin 3x = sin (90-3x) and use the general formula, \(3x=n\pi + (-1)^n (90-3x)\)

OpenStudy (anonymous):

that is the solution, not the proof though. So i don't think I gave any help into solving that question. note: \(n \in N\)

OpenStudy (anonymous):

u wanna calculate the value of x ?

OpenStudy (anonymous):

just find the vale of "X"

OpenStudy (anonymous):

sin 45 = cos 45 sin(3*15) = cos(3*15) so x=15 thus sin3x=cos3x your welcome

OpenStudy (anonymous):

thanks